A population data set has a mean of 81 and a standard deviation of 5.2. Find the z-scores for each of the following. Round to the nearest hundredth.

(a) x = 74

(b) x = 85
Answer #1

Solution :

(a)

x = 74

z = (x - ) / = (74 - 81) / 5.2 = -1.35

(b)

x = 85

z = (85 - 81) / 5.2 = 0.77